# An Efficient and Adaptive Watermark Detection System with Tile-based Error Correction

### v1

@@ 4.1 Pre-training the tile-based watermark extractor
At the start of the pre-training, we partition each training image into a regular grid of axis–aligned rectangles of size $l\times l$  px and randomly sample one grid cell $x_{t}\in R^{l\times l\times 3}$. Following the approach used in ReDMark [2], the encoder $H_{E}$ takes the tile as input along with a $N$-bit binary message $m\in\{0,1\}^{N}$ and then produces a residual image tile $\delta\in\mathbb{R}^{l\times l\times 3}$, that is further multiplied by a scalar hyperparameter $\alpha$ to derive the watermarked image tile $x_{w}=x_{0}+\alpha\delta$. At every optimization step, we first sample a transform $T$ from a set $\mathcal{T}$ that includes image transformations such as jpeg, and then apply $T$ to the watermarked image tile $x_{w}$. Eventually, the decoder $H_{D}$ then predicts a soft $N$-bit message $m^{\prime}=H_{D}(T(x_{w}))$ based on the transformed watermarked image $T(x_{w})$ as input. We set the standard binary cross-entropy loss as the message loss,
@@
where $\sigma(\cdot)$ denotes the sigmoid.
To incorporate the Reed–Solomon (RS) correction mechanism introduced later in §4.3, we introduce an RS-aware loss that penalizes only bit errors exceeding the correction capacity. Let $t$ denote the maximum number of symbol errors that an $(n,k)$ RS code can correct, and define the error indicator $e_{i}=\mathbbm{1}\!\left[\operatorname{sign}(m_{i}^{\prime})\neq m_{i}\right]$. To make sure errors within the RS capacity ($E\leq t$) incur no additional cost while uncorrectable errors are quadratically penalized, we formulated our RS-aware loss $\mathcal{L}_{\text{RS}}$ as $\max(0,\;E-t)]^{2}$ where $E=\sum_{i=1}^{k}e_{i}$. Thus, the overall training objective becomes
$\mathcal{L}=\mathcal{L}_{m}+\lambda\,\mathcal{L}_{\text{RS}},$
where $\lambda$ balances the standard bit error loss and the RS-aware term and is set to $\lambda=1$ in our experiments.
@@ 4.2 Fine-tuning the tile-based LDM decoder
[…]
Thus, the total objective is
$\mathcal{L}=\ \mathcal{L}_{\text{m}}\;+\;\lambda_{\text{i}}\;\mathcal{L}_{\text{i}}$
, with $\lambda_{\text{i}}=2.0$ unless stated otherwise. We fine-tune $\mathcal{D}_{m}$ for $100$ AdamW [21] iterations (batch size $4$) with 20 warm-up iterations to $10^{-4}$ followed by decay to $10^{-6}$.
@@ 4.3 Detection with Reed–Solomon Correction
During the watermark detection illustrated in Fig. 3(c), each candidate image is cropped and resized to $256\times 256$. The image is then partitioned into an $l\times l$ grid, where $l$ is either specified directly or predicted by the tile-size predictor described in Appendix B.2. From this grid, a single tile suffices for detection: the selected tile is passed through the decoder $H_{D}$, which outputs a raw bitstream $m^{\prime}$. Because the tile-based strategy processes only a portion of the image, the decoded bitstream may contain more errors than in the full-image Stable Signature model. To address this, Reed–Solomon (RS) error correction is applied to $m^{\prime}$ to obtain the final watermark message $c_{s}$, which is compared against the ground-truth $g$.
RS correction operates at the symbol level, with the maximum number of correctable errors given by $t=\left\lfloor\tfrac{m_{c}}{2}\right\rfloor$, where $m_{c}$ denotes the number of redundant correction symbols. In practice, we set the symbol size to 8 bits and use $m_{c}=2$, thereby enabling correction of up to one erroneous symbol. A systematic Berlekamp–Welch (B–W) decoder is then employed to reconstruct the original message from the noisy bitstream. Finally, the RS-corrected output is compared with the predetermined key for robust statistical verification. We show the full details of RS error correction in Appendix A.
@@ 5 Watermark Pipeline Optimization
In this section, we first present the image watermark detection pipeline, followed by the Adaptive Streams Allocation algorithm, which jointly optimizes stream allocation and batch partitioning, together with co-optimization of RS correction.
@@ 5.1 Image Watermark Detection Pipeline
Figure 4: Comparison of CUDA stream allocation strategies in the QRMark image watermark detection pipeline. (a) Adaptive CUDA Stream Allocation dynamically distributes streams across preprocessing, tiling, and decoding stages based on measured GPU utilization, thereby enabling greater stage overlap and reducing idle time. (b) Single Stream Strategy executes all stages strictly sequentially, often resulting in noticeable pipeline stalls during CPU-side Reed-Solomon correction.
The preprocessing stage consists of two substages: data loading and image transformation. As summarized in Table 1, QRMark applies a sequence of image processing operations to Stable Signature. The data loading substage transforms the raw image into a tensor and transfers it to the GPU. The transformation substage then performs CenterCrop (256), Resize (256), and Normalization. In the tiling stage, possible tiling strategy includes: random, random_grid, and fixed. We use random_grid as our default tiling strategy because it demonstrates higher robustness. Each watermarked image is divided into a regular grid composed of axis-aligned rectangular regions, each of size $n\times n$ pixels. A single grid cell is then randomly selected as the input for the decoding stage. In the decoding stage, the selected image tensor is processed by a pretrained extractor to recover the raw message bits. The raw message is then passed through our RS error correction module to obtain the final watermark. To mitigate the bottleneck introduced by RS correction, we employ an input queue together with a CPU-based thread pool. In addition, as described in §5.3, we introduce a caching mechanism to further reduce computational overhead. The CPU thread pool scales nearly linearly with the thread count $t$; in practice, we set $t=32$, which is sufficient for most workloads.
@@
Output: Streams per stage $s[1..K]$, mini-batches per stage $m[1..K]$
@@ 29     $\text{stall}\leftarrow\text{stall}+1$;
/* Step 3: Mini-batch leveling */
@@ 36     $m[k]\leftarrow\min(m_{\text{unit}},2m[k])$;
Algorithm 1 Adaptive Streams Allocation.
5.2 Adaptive Pipeline-level Stream Allocation
Inspired by the minibatch design in PipeDream [24], we propose a multi-stream allocation strategy in Fig. 4 to mitigate preprocessing and RS correction bottlenecks in the detection pipeline. Our primary design goal is to equalize the per-minibatch execution time across stages, thereby maximizing effective overlap and minimizing pipeline bubbles. Concretely, we partition GPU streams into three groups, each mapped to a pipeline stage. If stage $i$ receives $s_{i}$ streams, the global batch $B_{\text{cap}}$ is evenly divided into $s_{i}$ minibatches of size $B_{\text{cap}}/s_{i}$, and each minibatch is executed on an assigned CUDA stream. The number of streams per stage is chosen adaptively based on warm-up profiling that estimates each stage’s baseline time $t[i]$ and per-sample memory footprint $u[i]$. We allocate more streams to slower stages (large $t[i]$) to contract their step time, subject to a device-memory feasibility constraint $\sum_{i}s_{i}\cdot m_{i}\cdot u[i]\leq M_{\mathrm{cap}}$, where $m_{i}$ is the stage’s mini-batch size. We also enforce a global stream cap $\sum_{i}s_{i}\leq P$, trading off finer-grained balancing against launch costs.
As detailed in Algorithm 1, the allocation strategy begins with a warm-up phase to estimate each stage’s execution time and per-sample memory footprint (Line 1). It initializes the configuration by assigning one stream per stage and selecting the largest uniform mini-batch that fits the memory budget (Lines 2–4). It then computes the initial bottleneck latency $J^{\star}$ and resets the stall counter (Line 5). The algorithm enters the adaptive search loop (Line 8). Within each iteration, it evaluates stream augmentation: for every stage $k$, it tentatively increases $s[k]$ by one, and records the candidate configuration that yields the largest reduction in bottleneck latency (Lines 10–15). After scanning all stages, the algorithm decides whether to accept this best modification: if the improvement exceeds the threshold $\epsilon$, it updates $(s,m)$, recomputes $J^{\star}$, and resets the stall counter; otherwise, the stall counter is incremented (Lines 16–20). The search terminates once no further gains are observed or the stall counter reaches its cap. After the search halts, the algorithm performs mini-batch leveling: it derives a unit micro-batch $m_{\text{unit}}$ from the total stream count, and for stages substantially faster than the bottleneck, it increases their mini-batch size up to this bound (Lines 22–25). Finally, the algorithm returns the per-stage stream allocation and mini-batch sizes (Line 26).
@@ 5.3 Co-Optimizations for RS Correction
While adaptive pipeline-level stream allocation alleviates GPU-side contention, the detection pipeline can still suffer from stalls introduced by the Reed–Solomon (RS) correction stage. To address this, we apply optimizations at both the system and algorithmic levels. At the system level, we decouple RS correction from the GPU pipeline by introducing an input queue and a CPU thread pool. In §4.3, after the raw watermark messages $m^{\prime}$ are decoded, $m^{\prime}$ are dispatched to idle CPU threads for correction, and the final corrected outputs $c_{s}$ are collected asynchronously. This design prevents device-to-host transfers and CPU computation from stalling GPU progress. At the algorithm level, we observe that the embedded message sets are limited and detection accuracy is usually above 95%, leading to frequent recurrence of raw messages $m^{\prime}$. Inspired by that, we propose to maintain a codebook $cb$ that maps each $m^{\prime}$ to its corrected output $c_{s}$, together with a counter $c$ that tracks the number of images processed since its last access. When $m^{\prime}$ reappears, the cached result is directly reused, bypassing the costly correction step. The combination of system-level decoupling and algorithmic caching substantially reduces RS overhead while preserving correction fidelity.
@@ 6 Resource-aware Scheduling Optimization
In this section, we will demonstrate the inter-batch interleaving to overlap CPU-side data preparation and a resource-aware mini-batch schedule algorithm to balance workloads among streams and improve overall throughput.
@@ 6.1 Inter-Batch Workload Interleaving
The preprocessing stage in QRMark comprises two substages: CPU-side data preparation and GPU-side preprocessing. Inspired by the interleaving strategy in RAP [40], we adopt a similar approach that overlaps CPU preparation of the next batch with concurrent GPU kernel execution of the current batch, thereby achieving a more balanced overall schedule and reducing end-to-end latency.
@@
6.2 Resource-aware mini-batch Schedule
Building upon our adaptive CUDA stream allocation strategy and our interleaving strategy introduced in previous sections, we further propose a scheduling method to balance utilization at the mini-batch level. While interleaving hides preparation latency and decouples batch boundaries, the proposed scheduling method smooths per-stage load across streams to reduce pipeline bubbles during preprocessing, tiling, and decoding. In some scenarios, tiling and decoding involve strong data dependencies but small kernel footprints.
Input: Image set II; warm-up stats $\Pi$; pipeline mode; baseline batch $b_{0}$; stream budget $P$; balance slack $\lambda$; memory cap $M_{\mathrm{cap}}$; minimum mini-batch $b_{\min}$
Output: Schedule $S=\langle S_{1},\dots,S_{P}\rangle$; mini-batch size $m_{\text{unit}}$

### v2

@@ 4.1 Pre-training Tiled Watermark Extractor
In Fig. 3(a), the pre-training pipeline consists of a convolutional encoder and a convolutional extractor. The encoder learns to embed imperceptible perturbations into the cover image, while the extractor is simultaneously trained to recover the embedded message from the watermarked image. Both encoder and decoder are optimized by the HiDDeN [zhu2018hidden]. At the start of the pre-training, we partition each training image into a regular grid of axis–aligned rectangles of size $l\times l$  px and randomly sample one grid cell $x_{t}\in R^{l\times l\times 3}$. Following the approach used in ReDMark [redmark], the encoder $H_{E}$ takes the tile as input along with a $N$-bit binary message $m\in\{0,1\}^{N}$ and then produces a residual image tile $\delta\in\mathbb{R}^{l\times l\times 3}$, that is further multiplied by a scalar hyperparameter $\alpha$ to derive the watermarked image tile $x_{w}=x_{0}+\alpha\delta$. At every optimization step, we first sample a transform $T$ from a set $\mathcal{T}$ that includes image transformations such as jpeg, and then apply $T$ to the watermarked image tile $x_{w}$. Eventually, the decoder $H_{D}$ then predicts a soft $N$-bit message $m^{\prime}=H_{D}(T(x_{w}))$ based on the transformed watermarked image $T(x_{w})$ as input. We set the standard binary cross-entropy loss as the message loss,
@@
where $\sigma(\cdot)$ denotes the sigmoid. To incorporate the Reed-Solomon (RS) correction mechanism introduced later in §4.3, we introduce an RS-aware loss that penalizes only bit errors exceeding the correction capacity. Let $t$ denote the maximum number of symbol errors that an $(n,k)$ RS code can correct, and define the error indicator $e_{i}=\mathbbm{1}\!\left[\operatorname{sign}(m_{i}^{\prime})\neq m_{i}\right]$. To make sure errors within the RS capacity ($E\leq t$) incur no additional cost while uncorrectable errors are quadratically penalized, we formulated our RS-aware loss $\mathcal{L}_{\text{RS}}$ as $\max(0,\;E-t)]^{2}$ where $E=\sum_{i=1}^{k}e_{i}$. The overall training objective becomes
$\min\mathcal{L}=\mathcal{L}_{m}+\lambda\,\mathcal{L}_{\text{RS}},$
where $\lambda$ balances the standard bit error loss and the RS-aware term and is set to $\lambda=1$ in our experiments according to the settings in Stable Signature.
@@ 4.2 Fine-tuning the Tile-based LDM Decoder
The fine-tuning strategy closely mirrors that of Stable Signature, while it is adapted to our tile-based decoder. Concretely, we fine-tune the latent-diffusion decoder $\mathcal{D}_{m}$ such that the watermarked image it reconstructs encodes an $N$-bit watermark that can be reliably recovered by the tile extractor $H_{D}$ from the pre-training phase. As shown in Fig. 3(b), instead of explicitly embedding the fixed message $m$ in the image like what we did in §4.1, we first apply the RS encoding scheme (details in §4.3) to generate a binary signature $m_{s}$, with a bit length consistent with the pre-training stage, the $m_{s}$ consist of original message $m$ and correction symbols $m_{c}$ used for the RS correction purpose later in the watermark detection phase. For a training image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, we first obtain its latent $\mathbf{z}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{h\times w\times c},h=\tfrac{H}{f},w=\tfrac{W}{f}$ with the LDM encoder $\mathcal{E}$ where $f$ is a power-of-two down-sampling factor, $c$ is the dimension of channels for the latent code. Then, the decoder $\mathcal{D}_{m}$ transforms $\mathbf{z}$ into a watermarked image $\mathbf{x}^{\prime}=\mathcal{D}_{m}(\mathbf{z}),$ which is subsequently partitioned into non-overlapping $l\times l$ tiles. We randomly select one of the partitioned tiles and pass it to the extractor $H_{D}$ to predict the message $\mathbf{m}^{\prime}=H_{D}(\mathbf{x}^{\prime})$. To constrain distortion, we also reconstruct an unmarked image $\mathbf{x}_{o}^{\prime}=\mathcal{D}(\mathbf{z})$ with the frozen decoder $\mathcal{D}$ from the original latent diffusion model, and calculate the Watson-VGG perceptual loss [czolbe2021lossfunctiongenerativeneural] $\mathcal{L}_{\text{i}}=\operatorname{WatsonVGG}(\mathbf{x}^{\prime},\mathbf{x}_{o}^{\prime})$. Thus, the total objective is
$\mathcal{L}=\ \mathcal{L}_{\text{m}}\;+\;\lambda_{\text{i}}\;\mathcal{L}_{\text{i}},$
with $\lambda_{\text{i}}=2.0$ unless stated otherwise. We fine-tune $\mathcal{D}_{m}$ for $100$ AdamW [loshchilov2019decoupledweightdecayregularization] iterations (batch size $4$) with 20 warm-up iterations to $10^{-4}$ followed by decay to $10^{-6}$. As the parameters here are the same as the best setting in Stable Signature.
@@ 4.3 Detection with Reed–Solomon Correction
During the watermark detection illustrated in Fig. 3(c), each candidate image is cropped and resized to $256\times 256$. The image is then partitioned into an $l\times l$ grid, where $l$ is either specified directly or predicted by the tile-size predictor described in Appendix A.2. From this grid, a single tile suffices for detection: the selected tile is passed through the decoder $H_{D}$, which outputs a raw bitstream $m^{\prime}$. Because the tile-based strategy processes only a portion of the image, the decoded bitstream may contain more errors than in the full-image Stable Signature model. To address this, Reed–Solomon (RS) error correction is applied to $m^{\prime}$ to obtain the final watermark message $c_{s}$, which is compared against the ground-truth $g$.
RS correction operates at the symbol level, with the maximum number of correctable errors given by $t=\left\lfloor\tfrac{m_{c}}{2}\right\rfloor$, where $m_{c}$ denotes the number of redundant correction symbols. In practice, we set the symbol size to 8 bits and use $m_{c}=2$, thereby enabling correction of up to one erroneous symbol. A systematic Berlekamp-Welch (B-W) decoder is then employed to reconstruct the original message from the noisy bitstream. Empirically, we observe that our training and fine-tuning pipeline yields high per-tile extraction accuracy (>>90%) for most tile size settings. As a result, correcting a single erroneous symbol is sufficient to recover the small residual errors introduced during decoding and prevents further accuracy degradation in practice. Finally, the RS-corrected output is compared with the predetermined key for robust statistical verification. We provide full details of the RS error-correction procedure in Appendix B.
@@ 5 Watermark Pipeline Optimization
In this section, we first present the image watermark detection pipeline, followed by the adaptive multi-channel horizontal fusion algorithm, which jointly optimizes stream allocation and batch partitioning, together with co-optimization of RS correction.
@@ 5.1 Image Watermark Detection Pipeline
Figure 4: Illustration of optimizations applied in QRMark with multi-channel horizontal fusion and scheduling. Input images are preprocessed and partitioned into tiles on the GPU, followed by multi-channel decoding coordinated by adaptive resource management. Decoded messages are passed through a decoupled FIFO to the CPU-side Reed–Solomon correction for final watermark recovery.
Figure 5: Comparison of multi-channel horizontal fusion strategies in the QRMark image watermark detection pipeline. (a) multi-channel horizontal fusion dynamically distributes streams across preprocessing, tiling, and decoding stages based on measured GPU utilization, thereby enabling greater stage overlap and reducing idle time. (b) Single Stream Strategy executes all stages strictly sequentially, often resulting in noticeable pipeline stalls during CPU-side RS correction.
We show the optimizations applied to the detection pipeline in Fig. 4. The preprocessing stage consists of two substages: data loading and image transformation. As summarized in Table 1, QRMark applies a sequence of image processing operations to Stable Signature. The data loading substage transforms the raw image into a tensor and transfers it to the GPU. The transformation substage then performs CenterCrop (256), Resize (256), and Normalization. In the tiling stage, possible tiling strategy includes: random, random_grid, and fixed. We use random_grid as our default tiling strategy because it demonstrates higher robustness. Each watermarked image is divided into a regular grid composed of axis-aligned rectangular regions, each of size $n\times n$ pixels. A single grid cell is then randomly selected as the input for the decoding stage. In the decoding stage, the selected image tensor is processed by a pretrained extractor to recover the raw message bits. The raw message is then passed through our RS error correction module to obtain the final watermark. To mitigate the bottleneck introduced by RS correction, we employ an input queue together with a CPU-based thread pool. In addition, as described in §5.3, we introduce a caching mechanism to further reduce computational overhead. The CPU thread pool scales nearly linearly with the thread count $t$; in practice, we set $t=32$, which is sufficient for most workloads.
@@
Output: Streams per stage $s[1..K]$, micro-batches per stage $m[1..K]$
@@ 29     $\text{stall}\leftarrow\text{stall}+1$;
/* Step 3: micro-batch leveling */
@@ 36     $m[k]\leftarrow\min(m_{\text{unit}},2m[k])$;
Algorithm 1 Adaptive Multi-channel Horizontal Fusion.
5.2 Adaptive Multi-channel horizontal fusion
Inspired by the minibatch design in PipeDream [narayanan2019pipedream], we propose a multi-channel horizontal fusion strategy in Fig. 5 to mitigate preprocessing and RS correction bottlenecks in the detection pipeline. Our primary design goal is to equalize the per-minibatch execution time across stages, thereby maximizing effective overlap and minimizing pipeline bubbles. Concretely, we partition GPU streams into three groups, each mapped to a pipeline stage. If stage $i$ receives $s_{i}$ streams, the global batch $B_{\text{cap}}$ is evenly divided into $s_{i}$ minibatches of size $B_{\text{cap}}/s_{i}$, and each minibatch is executed on an assigned CUDA stream. The number of streams per stage is chosen adaptively based on warm-up profiling that estimates each stage’s baseline time $t[i]$ and per-sample memory footprint $u[i]$. We allocate more streams to slower stages (large $t[i]$) to contract their step time, subject to a device-memory feasibility constraint $\sum_{i}s_{i}\cdot m_{i}\cdot u[i]\leq M_{\mathrm{cap}}$, where $m_{i}$ is the stage’s micro-batch size. We also enforce a global stream cap $\sum_{i}s_{i}\leq P$, trading off finer-grained balancing against launch costs.
As detailed in Algorithm 1, the horizontal fusion strategy begins with a warm-up phase to estimate each stage’s execution time and per-sample memory footprint (Line 1). It initializes the configuration by assigning one stream per stage and selecting the largest uniform micro-batch that fits the memory budget (Lines 2–4). It then computes the initial bottleneck latency $J^{\star}$ and resets the stall counter (Line 5). The algorithm enters the adaptive search loop (Line 8). Within each iteration, it evaluates stream augmentation: for every stage $k$, it tentatively increases $s[k]$ by one, and records the candidate configuration that yields the largest reduction in bottleneck latency (Lines 10–15). After scanning all stages, the algorithm decides whether to accept this best modification: if the improvement exceeds the threshold $\epsilon$, it updates $(s,m)$, recomputes $J^{\star}$, and resets the stall counter; otherwise, the stall counter is incremented (Lines 16–20). The search terminates once no further gains are observed or the stall counter reaches its cap. After the search halts, the algorithm performs micro-batch leveling: it derives a unit micro-batch $m_{\text{unit}}$ from the total stream count, and for stages substantially faster than the bottleneck, it increases their micro-batch size up to this bound (Lines 22–25). Finally, the algorithm returns the per-stage stream allocation and micro-batch sizes (Line 26).
@@ 5.3 Co-Optimizations for RS Correction
While adaptive multi-channel horizontal fusion alleviates GPU-side contention, the detection pipeline can still suffer from stalls introduced by the Reed–Solomon (RS) correction stage. To address this, we apply optimizations at both the system and algorithmic levels. At the system level, we decouple RS correction from the GPU pipeline by introducing an input queue and a CPU thread pool. In §4.3, after the raw watermark messages $m^{\prime}$ are decoded, $m^{\prime}$ are dispatched to idle CPU threads for correction, and the final corrected outputs $c_{s}$ are collected asynchronously. This design prevents device-to-host transfers and CPU computation from stalling GPU progress. At the algorithm level, we observe that the embedded message sets are limited and detection accuracy is usually above 95%, leading to frequent recurrence of raw messages $m^{\prime}$. Inspired by that, we propose to maintain a codebook $cb$ that maps each $m^{\prime}$ to its corrected output $c_{s}$, together with a counter $c$ that tracks the number of images processed since its last access. When $m^{\prime}$ reappears, the cached result is directly reused, bypassing the costly correction step. The combination of system-level decoupling and algorithmic caching substantially reduces RS overhead while preserving correction fidelity.
@@ 6 Resource-aware Scheduling Optimization
In this section, we will demonstrate the inter-batch interleaving to overlap CPU-side data preparation and a resource-aware micro-batch schedule algorithm to balance workloads among streams and improve overall throughput.
@@ 6.1 Inter-Batch Workload Interleaving
The preprocessing stage in QRMark comprises two substages: CPU-side data preparation and GPU-side preprocessing. Inspired by the interleaving strategy in RAP [wang2024rap], we adopt a similar approach that overlaps CPU preparation of the next batch with concurrent GPU kernel execution of the current batch, thereby achieving a more balanced overall schedule and reducing end-to-end latency.
@@
6.2 Resource-aware micro-batch Schedule
Building upon our adaptive multi-channel horizontal fusion strategy and our interleaving strategy introduced in previous sections, we further propose a scheduling method to balance utilization at the micro-batch level. While interleaving hides preparation latency and decouples batch boundaries, the proposed scheduling method smooths per-stage load across streams to reduce pipeline bubbles during preprocessing, tiling, and decoding. In some scenarios, tiling and decoding involve strong data dependencies but small kernel footprints.
@@ 6.2 Resource-aware micro-batch Schedule
Input: Image set II; warm-up stats $\Pi$; pipeline mode; baseline batch $b_{0}$; stream budget $P$; balance slack $\lambda$; memory cap $M_{\mathrm{cap}}$; minimum micro-batch $b_{\min}$
Output: Schedule $S=\langle S_{1},\dots,S_{P}\rangle$; micro-batch size $m_{\text{unit}}$
2
31ex